Sara is considering two offers to purchase a car. Offer A is for $35000 consisting of an immediate $10,000 down payment with the $ 25,000 balance payable 1 year later. Offer B is for $33,500 made up of a $5,500 down payment and the $ 28,000 balance payable in 6 months. a. If money can earn 4% what is the current economic value of each offer? b. Other things being equal, which offer should Sara accept? What is the economic advantage of the preferred offer over the other offer?
1
Expert's answer
2020-06-02T18:58:00-0400

a)

Advance payments are paid at the beginning of the period, therefore, they will not give a profit.

Version A.

Define the profit from the use of money at 4% per annum.



"0.04\times25000=1000"

Then the total purchase price:



"35000-1000=34000"

Version B.

Define the profit from the use of money at 4% per annum.


"\frac{0.04\times28000}{2}=560"


Then the total purchase price:



"33500-560=32940"

b. Sara will accept version B.

Since we are talking about goods with similar consumer qualities, in this case the main one is the price paid for the car after taking into account all the possible benefits of using the money.